A silver-silver bromide wire is sometimes used as a reference electrode when Cl- -free conditions must be maintained. Write the cell notation and calculate the potential (at 30°C) of an Ag/AgBr reference electrode if it is immersed in a 3.5M solution of KBr. (Use concentrations, not activities in your calculation.)
Answer #1

The half cell notation for the Ag/AgBr electrode is

Br-(3.5M)/AgBr(s)/Ag(s)

Cell potential can be calculated using Nernst equation

E= E0-2.303RT{log[(Red)/(Oxd)]}/nF,

E= E0-2.303RT{log[Br-]}/nF

R = 8.314 j/K, T = 273+30 = 303K, n = 1 for Ag/AgBr electrode, F = 96500C (faraday constant), E0 = 0.0713 V (ref. Chem. Educ., 1969, 46 (11), p 741)

sustituting these values into Nernst equation

E = 0.0713-0.0601log(3.5)

E = 0.0385 V